The manufacturer of a new drug claims that it will lower

blood pressure by 13 points on the average. When the drug was

administered to 6 patients, the following drops in blood pressure

were registered: 12, 8, 15, 9, 10 and 16. Is the claim sustained at

the 0.05 level of significance.


1
Expert's answer
2022-05-19T17:08:27-0400

Mean "\mu" = "\dfrac{12+8+15+9+10+16}{6}=\dfrac{35}{3}"

Variance of population 


"s^2=\dfrac{\Sigma(x_i-\bar{x})^2}{n-1}=\dfrac{1}{6-1}((12-\dfrac{35}{3})^2"


"+(8-\dfrac{35}{3})^2+(15-\dfrac{35}{3})^2+(9-\dfrac{35}{3})^2"

"+(10-\dfrac{35}{3})^2+(16-\dfrac{35}{3})^2)=\dfrac{32}{3}"

"s=\sqrt{s^2}=\sqrt{\dfrac{32}{3}}\approx3.266"


The following null and alternative hypotheses need to be tested:

"H_0:\mu=13"

"H_a:\mu\not=13"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.05," "df=n-1=5" degrees of freedom, and the critical value for a two-tailed test is "t_c =2.570543."The rejection region for this two-tailed test is "R = \{t:|t|> 2.570543\}."

The t-statistic is computed as follows:


"t=\dfrac{11.667-13}{3.266/\sqrt{6}}=-0.9997"

Since it is observed that "|t| = 0.9997<2.570543=t_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value for two-tailed "df=5" degrees of freedom, "t=-0.9997" is "p=0.363349," and since "p=0.363349>0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu" is different than 13, at the "\alpha = 0.05" significance level.